OpenStudy (prince1342):

Point A is located at (0, 4) and point B is located at (−2, −3). Find the x value for the point that is 1 over 4 the distance from point A to point B.

OpenStudy (prince1342):

−1.5 −2 −0.5 −1

OpenStudy (prince1342):

@FaiqRaees Help please, its my last question

OpenStudy (faiqraees):

\[distance = \sqrt{(x2-x1)^2 + (y2-y1)^2}\] first find the distance between A and B

OpenStudy (prince1342):

I got 53

OpenStudy (retireed):

Brain misfire.

OpenStudy (faiqraees):

oh sorry it can just be calculated by 0 +(-2)/4 which gives us -1/2 add that to the x coordinate of A x= 0 + (-1/2) x= -1/2

OpenStudy (retireed):

A much better answer. My first attempt I was thinking of quadrant 4 instead of quadrant 3.

OpenStudy (retireed):



OpenStudy (prince1342):

Hmm so with that said, what wwould it make it be?

OpenStudy (prince1342):

Sorry my reply is late, I had to go to my violin lessons

OpenStudy (prince1342):

-1/2 would be -0.5 basically

OpenStudy (retireed):